# Safe and Optimal N-Spacecraft Swarm Reconfiguration in Non-Keplerian Cislunar Orbits

### v1

@@ 5.2.2 BCR4BP $L_{2}$ South 3:1 Halo Orbit (synodic resonance)
The optimization results are presented in Fig. 16 in the same fashion to Fig. 15. First, as shown in Fig. 16(a), the shape of the QPRIT is remarkably different from the previous two cases. The perturbation of the absolute orbit leads to the highly eccentric invariant curve. The optimized transfer cost of the CVX, CVX-QPRIT, and SCP-QPRIT is $\Delta v_{tot}$ = 2.051 mm/s, 5.559 mm/s, and 5.637 mm/s, respectively. It is observed that the SCP-QPRIT solution stays along the surface of the torus (cf. Figs. 16(a) and 16(d)) before reducing $\varepsilon$. The CVX-QPRIT solution almost satisfies passive safety, where only slight modification is needed in the following SCP to ensure the constraint in Eq. 21b. Another key observation is the high sensitivity of the plan OCP solution in the LTC space. This is due to the rapid rotation of the invariance circle, where the slight change in the Cartesian states leads to a spike in the LTC when the change is orthogonal to the change of LTC over time. This sensitivity is suppressed by simply imposing the constraint in Eq. 21a. The terminal nonlinear error between the SCP-QPRIT solution under the linearized dynamics and the nonlinear dynamics of BCR4BP is 0.657m, which is sufficiently low for the open-loop guidance and control of this formation reconfiguration problem.
(a) TNW frame
(b) Control History (TNW frame)
@@ 5.3 N-spacecraft Swarm Reconfiguration in a Full-Ephemeris Dynamics Model
As the final case study, the proposed method is validated in a five-spacecraft swarm reconfiguration in the $L_{2}$ South 9:2 synodic NRHO under the full-ephemeris dynamics model, where exact periodic solutions are not attainable. The full-ephemeris model is based on the JPL SPICE toolkit  [^6] https://naif.jpl.nasa.gov/naif/toolkit.html and considers the point-mass model of the Sun-Earth-Moon system, where the initial epoch is set to 2025-01-01 00:00:00 UTC. Other perturbations, such as spherical harmonics or solar radiation pressure, are not considered. The validation process is illustrated in Fig. 17. Because the LTC (and an oscillatory mode) is well-defined only if the reference absolute orbit is perfectly periodic, an optimal relative guidance trajectory is generated within the simplified dynamics model, which permits periodic solutions. In this case study, the trajectory is optimized in the relative dynamics of CR3BP. In contrast, the chief orbit is separately obtained by converging a trajectory in the full-ephemeris model. In this case study, the 9-revolutions of NRHO is continued using multiple-shooting with the Newton-Raphson method from the CR3BP solution [48]. The full-ephemeris NRHO is shown in Fig. 10(a). The optimal relative control profile obtained in the first step is then propagated in the full-ephemeris model along with the chief orbit that is continued in the full-ephemeris model. Note that the same dynamics model that is used to continue the full-ephemeris NRHO is also used for the propagation of relative motion. By performing this validation process, the station-keeping of the absolute orbit is effectively neglected from both the chief and deputy trajectories. This enables the extraction of the error in relative dynamics models between the simplified dynamics model (CR3BP) and the full-ephemeris dynamics. Since the propagation period in this case study is up to a few revolutions, this can serve as a simple validation scheme of the relative orbital control based on the proposed method.
@@
Because of the primitiveness of the CR3BP, the dynamics modeling error causes a large deviation when the feed-forward control is propagated in the full-ephemeris dynamics model. An MPC framework is introduced to address this issue, where the nonconvex trajectory optimization is re-solved via the SCP for the remaining horizon at predefined epochs $t_{mpc}$. The corresponding time step indices $k_{mpc}$ are listed in Table 2. Throughout the transfer, a history of STM, control input matrix, and eigenvector $\boldsymbol{w}(t)$ is carried over (i.e., the LTC is not re-initialized). In this case study, navigation uncertainty is not explicitly considered, and each agent is assumed to have perfect knowledge, while more sophisticated event-based MPC may be considered using the state covariance information [10]. However, to account for the actuation error, unmodeled acceleration (e.g., differential in spherical harmonics and solar radiation pressure), a linear process noise is added to the dynamics propagation, modeled as $Q(t_{k+1}-t_{k})$, where $Q$ represents the noise matrix (cf. Table 2).
@@
Figure 18: 100 Monte-Carlo simulation of the swarm reconfiguration in the full-ephemeris dynamics model. Sky-blue trajectories are based on the Open-Loop (OL) control, whereas grey trajectories are based on the MPC scheme.
@@
$\mathcal{J}_{j}$ [mm/s] $\rho_{\text{TNW}}$ [m] $\dot{\rho}_{\text{TNW}}$ [mm/s] $h$ [m] $\alpha$ [m] $\beta$ [m] $\dot{h}$ [mm/s] $\dot{\alpha}$ [mm/s] $\dot{\beta}$ [mm/s]
@@
Figure 19: History of Inter-spacecraft separation in nominal case (upper/lower-bound) in CR3BP and the worst separation of each combinatorics in 100 Monte-Carlo simulations in the full-ephemeris model.
As a final note, it is observed through extensive experiments that the epochs that trigger the MPC scheme for trajectory replanning are tightly coupled with the full-ephemeris reference orbit. Particularly, slight changes in the full-ephemeris NRHO result in completely different behavior on the CR3BP’s LTC space, which poses an additional challenge in the practical deployment of a spacecraft swarm. Therefore, in future research, an integrated absolute-relative motion control should be addressed.
@@ 6 Conclusion
Safety of bounded relative motion in Restricted Multi-Body Problems (RMBPs) is one of the largest concerns in the practical deployment of spacecraft swarms in non-Keplerian cislunar orbits. This paper presents a novel passively safe relative reconfiguration strategy for RMBPs. The main contribution is the formulation of the optimal control problem in the Local Toroidal Coordinates (LTC), a time-varying frame that projects the relative motion onto the oscillatory eigenspace that enables the bounded quasi-periodic relative motion. In the LTC, safety constraints drawn from the quasi-periodic structure are significantly simplified and decentralized between each agent, enabling the large-scale swarm reconfiguration. To support this method, the equations of relative motion of the PRMBPs are re-derived in the TNW frame, a co-moving frame that aligns with the spacecraft’s velocity direction. TNW frame is advantageous to discuss the relative motion in highly elliptic orbits, such as Near Rectilinear Halo Orbits (NRHOs), as it slows the natural relative motion in the radial direction. Furthermore, geometric relationships of the bounded quasi-periodic relative motion in the RMBPs and the bounded periodic relative motion in the Restricted Two-Body Problem (R2BP) are analyzed through their local eigensystems. The proposed reconfiguration strategy is demonstrated in various cislunar NRHOs in the Circular Restricted Three-Body Problem, Elliptic Three-Body Problem, and Bicircular Restricted Four-Body Problem. Furthermore, the conversion of the guidance strategy in the full-ephemeris model is successfully performed via model predictive control. This paper builds a practical stepping stone for the deployment of large-scale spacecraft swarming operations in a highly perturbed dynamics in the cislunar space. Future research includes the integration of absolute orbit station-keeping and tracking schemes into the relative motion control and more streamlined swarm design optimization methods.
@@ 6.1.2 BCR4BP
6.2 Derivation of the Angular Velocoity and Acceleration of the TNW Frame to the Moon Synodic Frame
The derivation of the angular velocity of the TNW frame with respect to the Moon synodic frame $\boldsymbol{\omega}_{v/m}$ follows the procedure elaborated in Ref. [16]. The derivation first requires a time derivative of basis vectors of the TNW frame expressed in the Moon synodic frame: $\left\{[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}},[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}},[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}\right\}$.
@@ 6.2 Derivation of the Angular Velocoity and Acceleration of the TNW Frame to the Moon Synodic Frame
$[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}$ $=\dfrac{d}{dt}\left(\dfrac{[\dot{\boldsymbol{r}}]_{\mathcal{M}}}{v}\right)=\dfrac{1}{v}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}-\hat{\boldsymbol{\imath}}\cdot(\hat{\boldsymbol{\imath}}\cdot[\dot{\boldsymbol{r}}]_{\mathcal{M}}))\quad\left(\because\dot{v}=\dfrac{[\dot{\boldsymbol{r}}]_{\mathcal{M}}\cdot[\ddot{\boldsymbol{r}}]_{\mathcal{M}}}{v}=\hat{\boldsymbol{\imath}}\cdot[\ddot{\boldsymbol{r}}]_{\mathcal{M}}\right)$ (27a)
$=\dfrac{1}{v}(([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{\jmath}})\hat{\boldsymbol{\jmath}}+([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}})\hat{\boldsymbol{k}})$ (27b)
@@
$[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}$ $=\dfrac{1}{h^{2}}([\dot{\boldsymbol{h}}]_{\mathcal{M}}h-\boldsymbol{h}\dot{h})$ (28a)
$=\dfrac{1}{h}([\dot{\boldsymbol{h}}]_{\mathcal{M}}-\dot{h}{\hat{\boldsymbol{\jmath}}})$ (28b)
@@
From Eq. 28b to 28c, the following equality is used, noting that $[\dot{\boldsymbol{h}}]_{\mathcal{V}}=\dot{h}\hat{\boldsymbol{\jmath}}$:
$\dot{h}$ $=[\dot{\boldsymbol{h}}]_{\mathcal{V}}\cdot\hat{\boldsymbol{\jmath}}=[\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{\jmath}}-(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\hat{\boldsymbol{\jmath}}=[\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{\jmath}}.\quad\left(\because(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\hat{\boldsymbol{\jmath}}=\dfrac{1}{h}(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\boldsymbol{h}=0\right)$ (29a)
Also note that $[\dot{\boldsymbol{h}}]_{\mathcal{M}}=\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}}$. Finally, the time derivative of $\hat{\boldsymbol{k}}$ is computed as follows:
$[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}$ $=[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}\times\hat{\boldsymbol{\jmath}}+\hat{\boldsymbol{\imath}}\times[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}=-\dfrac{1}{v}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}})\hat{\boldsymbol{\imath}}-\dfrac{1}{h}([\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot{\hat{\boldsymbol{k}}}){\hat{\boldsymbol{\jmath}}}.$ (30)
The time derivative of unit vectors of the TNW frame expressed in the Moon synodic frame can also be expressed using $\boldsymbol{\omega}_{v/m}$ as follows:
@@
$\hat{\boldsymbol{\imath}}\times[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{\imath}})\hat{\boldsymbol{\imath}},\quad\hat{\boldsymbol{\jmath}}\times[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{\jmath}})\hat{\boldsymbol{\jmath}},\quad\hat{\boldsymbol{k}}\times[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{k}})\hat{\boldsymbol{k}}.$ (32)
@@
$\boldsymbol{\omega}_{v/m}=\dfrac{1}{2}\left(\hat{\boldsymbol{\imath}}\times[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}+\hat{\boldsymbol{\jmath}}\times[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}+\hat{\boldsymbol{k}}\times[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}\right)$ (33)
With Eqs. (27), (28), and (30), the angular velocity $\boldsymbol{\omega}_{v/m}$ is expressed in the TNW frame as follows:
@@
Each component of $[\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}$ is computed by simply taking the component-wise time derivative of the above expression, which results in
$\dot{\omega}^{x}_{v/m}$ $=-\dfrac{\dot{h}}{h^{2}}((\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}})+\dfrac{1}{h}\left(([\dot{\boldsymbol{r}}]_{\mathcal{M}}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}}+(\boldsymbol{r}\times[\dddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}}+(\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}\right)$ (35a)
$\dot{\omega}^{y}_{v/m}$ $=\dfrac{\dot{v}}{v^{2}}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}})-\frac{1}{v}([\dddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}}+[\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}})$ (35b)

### v2

@@ 5.2.2 BCR4BP $L_{2}$ South 3:1 Halo Orbit (synodic resonance)
The optimization results are presented in Fig. 16 in the same fashion to Fig. 15. First, as shown in Fig. 16(a), the shape of the QPRIT is remarkably different from the previous two cases. The perturbation of the reference orbit leads to the highly eccentric invariant curve. The optimized transfer cost of the CVX, CVX-QPRIT, and SCP-QPRIT is $\Delta v_{tot}$ = 2.051 mm/s, 5.559 mm/s, and 5.637 mm/s, respectively. It is observed that the SCP-QPRIT solution stays along the surface of the torus (cf. Figs. 16(a) and 16(d)) before reducing $\varepsilon$. The CVX-QPRIT solution almost satisfies passive safety, where only slight modification is needed in the following SCP to ensure the constraint in Eq. 24b. Another key observation is the high sensitivity of the plan OCP solution in the LTC space. This is due to the rapid rotation of the invariance circle, where the slight change in the Cartesian states leads to a spike in the LTC when the change is orthogonal to the change of LTC over time. This sensitivity is suppressed by simply imposing the constraint in Eq. 24a. The terminal nonlinear error between the SCP-QPRIT solution under the linearized dynamics and the nonlinear dynamics of BCR4BP is 0.657m, which is sufficiently low for the open-loop guidance and control of this formation reconfiguration problem.
(a) VNB
frame
(b) Control History (VNB
frame)
@@ 5.3 N-spacecraft Swarm Reconfiguration in a Full-Ephemeris Dynamics Model
As the final case study, the proposed method is validated in a five-spacecraft swarm reconfiguration in the $L_{2}$ South 9:2 synodic NRHO under the full-ephemeris dynamics model, where exact periodic solutions are not attainable. The full-ephemeris model is based on the JPL SPICE toolkit  [^6] https://naif.jpl.nasa.gov/naif/toolkit.html and considers the point-mass model of the Sun-Earth-Moon system, where the initial epoch is set to 2025-01-01 00:00:00 UTC. Other perturbations, such as spherical harmonics or solar radiation pressure, are not considered. The validation process is illustrated in Fig. 17. Because the LTC (and an oscillatory mode) is well-defined only if the reference orbit is perfectly periodic, an optimal relative guidance trajectory is generated within the simplified dynamics model, which permits periodic solutions. In this case study, the trajectory is optimized in the relative dynamics of CR3BP. In contrast, the chief orbit is separately obtained by converging a trajectory in the full-ephemeris model. In this case study, the 9-revolutions of NRHO is continued using multiple-shooting with the Newton-Raphson method from the CR3BP solution [62]. The full-ephemeris NRHO is shown in Fig. 10(a). The optimal relative control profile obtained in the first step is then propagated in the full-ephemeris model along with the chief orbit that is continued in the full-ephemeris model. Note that the same dynamics model that is used to continue the full-ephemeris NRHO is also used for the propagation of relative motion. By performing this validation process, the station-keeping of the reference orbit around the Moon is effectively neglected from both the chief and deputy trajectories. This enables the extraction of the error in relative dynamics models between the simplified dynamics model (CR3BP) and the full-ephemeris dynamics. Since the propagation period in this case study is up to a few revolutions, this can serve as a simple validation scheme of the relative orbital control based on the proposed method.
@@
Because of the simplifying approximation of the CR3BP, the dynamics modeling error causes a large deviation when the feed-forward control is propagated in the full-ephemeris dynamics model. A shrinking-horizon MPC framework is introduced to address this issue, where the nonconvex trajectory optimization is re-solved via the SCP for the remaining horizon at predefined epochs $t_{\text{MPC}}$. The corresponding time step indices $k_{\text{MPC}}$ are listed in Table 2, where the OCP is resolved twice during the transfer at $k_{\text{MPC}}=$ 13 and 25. Throughout the transfer, a history of STM, control input matrix, and eigenvector $\boldsymbol{w}(t)$ is carried over (i.e., the LTC is not re-initialized). In this case study, navigation uncertainty is not explicitly considered, and each agent is assumed to have perfect knowledge, while more sophisticated event-based MPC may be considered using the state covariance information [13]. However, to account for the actuation error, unmodeled acceleration (e.g., differential in spherical harmonics and solar radiation pressure), a linear process noise is added to the dynamics propagation, modeled as $Q(t_{k+1}-t_{k})$, where $Q$ represents the noise matrix (cf. Table 2).
@@
Figure 18: Trajectories from 100 Monte-Carlo simulations of swarm reconfiguration in the full‑ephemeris dynamics model under Open‑Loop (OL) control and MPC.
@@
$\mathcal{J}_{j}$ [mm/s] $\rho_{\text{VNB}}$ [m] $\dot{\rho}_{\text{VNB}}$ [mm/s] $h$ [m] $\alpha$ [m] $\beta$ [m] $\dot{h}$ [mm/s] $\dot{\alpha}$ [mm/s] $\dot{\beta}$ [mm/s]
@@
Figure 19: Comparison of inter‑spacecraft separation: the nominal bounds in the SCP solution and the worst‑case separations between each agent from 100 Monte‑Carlo simulations in the full-ephemeris model (Open-Loop (OL) and MPC trajectories).
As a final note, it is observed from the extensive testing that MPC trigger epochs are highly sensitive to the chosen full‑ephemeris reference orbit. Even minor perturbations to the full‑ephemeris NRHO can produce different trajectories in the CR3BP LTC space, complicating swarm‑level implementation. Therefore, in future research, a concurrent control scheme of station-keeping of reference orbit and around the Moon and formation reconfiguration between spacecraft should be addressed.
@@ 6 Conclusion
Safety of bounded relative motion in Restricted Multi-Body Problems (RMBPs) is one of the largest concerns in the practical deployment of spacecraft swarms in non-Keplerian cislunar orbits. This paper presents a novel passively safe relative reconfiguration strategy for RMBPs. The main contribution is the formulation of the optimal control problem in the Local Toroidal Coordinates (LTC), a time-varying frame that projects the relative motion onto the oscillatory eigenspace that enables the bounded quasi-periodic relative motion. In the LTC, safety constraints drawn from the quasi-periodic structure are significantly simplified and decentralized between each agent, enabling the large-scale swarm reconfiguration. To support this method, relative motion of the PRMBPs is resolved in the Velocity/Normal/Binormal(VNB) frame, a co-moving frame that aligns with the spacecraft velocity. The VNB frame is advantageous to discuss the relative motion in highly elliptic orbits, such as Near Rectilinear Halo Orbits (NRHOs), as it slows the natural relative motion in the radial direction. Furthermore, geometric relationships of the bounded quasi-periodic relative motion in the RMBPs and the bounded periodic relative motion in the Perturbed Restricted Two-Body Problem (PR2BP) are analyzed through their local eigensystems. The proposed reconfiguration strategy is demonstrated in various cislunar NRHOs in the Circular Restricted Three-Body Problem, Elliptic Three-Body Problem, and Bicircular Restricted Four-Body Problem. Furthermore, the deployment of the guidance strategy in the full-ephemeris model is successfully performed through model predictive control. This paper builds a practical stepping stone for the deployment of large-scale spacecraft swarming operations in a highly perturbed dynamics in cislunar space. Future research includes the integration of station-keeping of a reference orbit and relative motion control, as well as a streamlined swarm design optimization method based on obaservability analysis.
@@ 6.1.2 BCR4BP
6.2 Derivation of the Angular Velocoity and Acceleration of the VNB Frame to the Moon Synodic Frame
The derivation of the angular velocity of the VNB frame with respect to the Moon synodic frame $\boldsymbol{\omega}_{v/m}$ follows the procedure elaborated in Ref. [19]. The derivation first requires a time derivative of basis vectors of the VNB frame expressed in the Moon synodic frame: $\left\{[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}},[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}},[\dot{\hat{\boldsymbol{k}}}$$_{\mathcal{M}}\right\}$.
@@ 6.2 Derivation of the Angular Velocoity and Acceleration of the VNB Frame to the Moon Synodic Frame
$[\dot{\hat{\boldsymbol{\imath}}}$$_{\mathcal{M}}$ $=\dfrac{d}{dt}\left(\dfrac{[\dot{\boldsymbol{r}}]_{\mathcal{M}}}{v}\right)=\dfrac{1}{v}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}-\hat{\boldsymbol{\imath}}\cdot(\hat{\boldsymbol{\imath}}\cdot[\dot{\boldsymbol{r}}]_{\mathcal{M}}))\quad\left(\because\dot{v}=\dfrac{[\dot{\boldsymbol{r}}]_{\mathcal{M}}\cdot[\ddot{\boldsymbol{r}}]_{\mathcal{M}}}{v}=\hat{\boldsymbol{\imath}}\cdot[\ddot{\boldsymbol{r}}]_{\mathcal{M}}\right)$ (30a)
[…]
@@
$[\dot{\hat{\boldsymbol{\jmath}}}$$_{\mathcal{M}}$ $=\dfrac{1}{h^{2}}([\dot{\boldsymbol{h}}]_{\mathcal{M}}h-\boldsymbol{h}\dot{h})$ (31a)
[…]
@@
From Eq. 31b to 31c, the following equality is used, noting that $\dot{\boldsymbol{h}}]_{\mathcal{V}}=\dot{h}\hat{\boldsymbol{\jmath}}$:
$\dot{h}$ $=[\dot{\boldsymbol{h}}]_{\mathcal{V}}\cdot\hat{\boldsymbol{\jmath}}=[\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{\jmath}}-(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\hat{\boldsymbol{\jmath}}=[\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{\jmath}}.\qquad\left(\because(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\hat{\boldsymbol{\jmath}}=\dfrac{1}{h}(\boldsymbol{\omega}_{v/m}\times\boldsymbol{h})\cdot\boldsymbol{h}=0\right)$ (32a)
Also note that $\dot{\boldsymbol{h}}]_{\mathcal{M}}=\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}}$. Finally, the time derivative of $\hat{\boldsymbol{k}}$ is computed as follows:
$[\dot{\hat{\boldsymbol{k}}}$$_{\mathcal{M}}$ $=[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}\times\hat{\boldsymbol{\jmath}}+\hat{\boldsymbol{\imath}}\times[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}=-\dfrac{1}{v}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}})\hat{\boldsymbol{\imath}}-\dfrac{1}{h}([\dot{\boldsymbol{h}}]_{\mathcal{M}}\cdot{\hat{\boldsymbol{k}}}){\hat{\boldsymbol{\jmath}}}$. (33)
The time derivative of unit vectors of the VNB frame expressed in the Moon synodic frame can also be expressed using $\boldsymbol{\omega}_{v/m}$ as follows:
@@
$\hat{\boldsymbol{\imath}}\times[\dot{\hat{\boldsymbol{\imath}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{\imath}})\hat{\boldsymbol{\imath}},\qquad\hat{\boldsymbol{\jmath}}\times[\dot{\hat{\boldsymbol{\jmath}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{\jmath}})\hat{\boldsymbol{\jmath}},\qquad\hat{\boldsymbol{k}}\times[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}=\boldsymbol{\omega}_{v/m}-(\boldsymbol{\omega}_{v/m}\cdot\hat{\boldsymbol{k}})\hat{\boldsymbol{k}}$. (35)
@@
[…]
With Eqs. (30), (31), and (33), the angular velocity $\boldsymbol{\omega}_{v/m}$ is expressed in the VNB frame as follows:
@@
Each component of $\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}$ is computed by simply taking the component-wise time derivative of the above expression, which results in
[…]